OpenStudy (bloomlocke367):

I just need to see if I'm doing this correctly.. question in the comments.

OpenStudy (bloomlocke367):

Find the sum of the finite series.\[\sum_{n=1}^{10}(3n-4)\] So do I just plug in 1, 2, 3, 4, 5, 6, 7, 8, 9, and 10 in for n and add them all together?

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (phi):

that will work, but there are shorter ways to do it.

OpenStudy (bloomlocke367):

what's the shorter way?

OpenStudy (bloomlocke367):

because right now, I have -1+2+5+8+11+14+17+20+23+24

OpenStudy (bloomlocke367):

=123

OpenStudy (bloomlocke367):

@phi

OpenStudy (bloomlocke367):

@ganeshie8

OpenStudy (bloomlocke367):

oops, not 24, 26

OpenStudy (bloomlocke367):

so it equals 125

OpenStudy (phi):

the short way is write the problem as \[ \sum_{n=1}^{10}(3n-4) = 3\sum_{n=1}^{10}n- \sum_{n=1}^{10} 4\] the last summation is the sum of ten 4's, which is 10*4= 40. make it negative because we will be subtracting this: -40 the first summation is the sum of the first 10 numbers there is a formula to find the sum of the first n numbers (starting with 1) it is n(n+1)/2 here n is 10, so the sum of the first 10 numbers is 10*11/2 = 5*11= 55 now multiply that by 3: 3*55= 165 total is 165-40= 125